OpenStudy (anonymous):

simplify. 8! over 6! =?

OpenStudy (anonymous):

8!/ 6! = (8 x 7 x 6!)/ 6! = 8 x 7 = 56

OpenStudy (across):

\[\frac{8!}{6!}=\frac{8*7*6*5*4*3*2*1}{6*5*4*3*2*1}=8*7\]

OpenStudy (anonymous):

oh i get it now! thank you:)

OpenStudy (anonymous):

welcome